# Realistic Market Impact Modeling for Reinforcement Learning Trading Environments

### 30 Mar 2026 v1

@@ Abstract
Reinforcement learning (RL) has shown promise for trading, yet most open-source backtesting environments assume negligible or fixed transaction costs, which can cause agents to learn undesirable trading behaviors that fail on real-world execution. We introduce a suite of three Gymnasium-compatible trading environments—MACE (Market-Adjusted Cost Execution) stock trading, margin trading, and portfolio optimization—that integrate nonlinear market impact models grounded in the Almgren–Chriss (AC) framework and the empirically validated square-root impact law. Each environment shares a unified infrastructure: pluggable cost models, permanent impact tracking with exponential decay, and comprehensive trade-level logging. We evaluate five DRL algorithms (A2C, PPO, DDPG, SAC, TD3) on the NASDAQ 100 universe, comparing a fixed 10 bps baseline cost model against the AC model with Optuna-optimized hyperparameters. Our results show that (i) the cost model materially affects both absolute performance and the relative ranking of algorithms across all three environments—in stock trading, optimized PPO achieves the best OOS result (20% return, Sharpe 1.06) under the baseline but drops to 15% under AC, while TD3 improves from 15% to 18%; in portfolio optimization, TD3 under AC achieves 32% (best overall) while the same algorithm under baseline yields only 26% (worst among agents); (ii) the AC model produces dramatically different trading behavior—non-optimized TD3 in stock trading sees daily costs drop from \$200k to \$8k when switching to AC, with turnover falling from 19% to 1%; (iii) hyperparameter optimization (HPO) is essential for constraining pathological trading—without hyperparameter optimization, agents exhibit monotonically increasing participation rates across epochs (e.g., SAC stock trading costs drop 82% with HPO); and (iv) algorithm choice interacts strongly with the cost model in environment-specific ways—in the margin environment, DDPG’s OOS Sharpe jumps from -2.1 to 0.3 under AC while SAC’s drops from -0.5 to -1.2. We release the full environment suite as an open-source extension to FinRL-Meta.
@@ III-C Reward Function
Figure 1: OOS return—MACE stock trading, all five agents under baseline vs. AC impact, optimized params. Black line: QQEW benchmark (19%).
@@ V-A MACE Stock Trading: Five-Agent Comparison
All configurations outperform the QQEW benchmark (19% OOS return) except non-optimized TD3 under AC. The best result is optimized PPO under the baseline cost model: 20% annualized return and 1.06 Sharpe, switching to the AC cost model reduces return to 15% (Sharpe 1.03), but trading costs drop 55%, average order POV from 0.18 to 0.14, max drawdown from -20% to -16%, turnover from 1.3% to 1.1%, and volatility from 19% to 15%—a substantially less risky portfolio. TD3 shows the opposite pattern: AC improves returns from 15% to 18% and Sharpe from 0.9 to 1.1, while turnover drops from 4% to 3%, costs fall 13%, and POV decreases from 0.36 to 0.20. DDPG also improves modestly under AC; A2C and SAC are slightly worse.
HPO dramatically reduces trading activity. Optimized SAC cuts turnover from 5% to 2% on AC and trading costs by 82% (POV from 0.46 to 0.18); DDPG shows a similar pattern. All optimized agents place heavy emphasis on cost reduction, sometimes at the expense of raw returns. The most extreme case is non-optimized TD3 under the baseline: average order POV of 1%, daily turnover of 19%, and average daily trading cost of \$200k. The same agent under AC achieves POV of 0.2%, turnover of 1%, and daily cost of \$8k—a 96% reduction that only moderately affects returns (13% vs. 11%) because lower volatility (17% to 12%) preserves Sharpe (0.8 to 0.9) (Fig. 2).
@@ V-B Margin Trading: Four-Agent Comparison
All agents underperform the QQEW benchmark OOS, although A2C and PPO outperform IS, suggesting overfitting that HPO could mitigate. The AC model’s effect varies significantly by agent. A2C returns 5.5% (baseline) and 6.5% (AC) OOS, with IS returns of 13% and 17.5% respectively—the IS–OOS gap is large but the AC model produces slightly better OOS performance. PPO achieves 15% OOS under the baseline but only 9% under AC, despite similar turnover ($\sim$3%) and 40% lower trading costs on AC. DDPG and SAC are the weakest performers, and the cost model makes a large difference in opposite directions: DDPG improves dramatically under AC (OOS Sharpe from -2.1 to 0.3, annualized return from -11% to 2%, max drawdown from -23% to -6%) with similar trading costs and POV, while SAC degrades under AC (OOS Sharpe from -0.5 to -1.2, return from -3% to -7%), with lower POV (0.26 to 0.15) but $\sim$10% higher trading costs (Fig. 4).
@@ V-C Portfolio Optimization: Five-Agent Comparison
All agents outperform the benchmark. TD3 exhibits the widest spread across cost models: with the AC model it achieves 32%—the highest OOS return of any configuration—while with the 10 bps baseline it returns 26%, the lowest among all agents. A2C obtains $\sim$30% under both models (slightly higher with AC); SAC reaches 30% (baseline) and 29% (AC); DDPG is comparable at 28% (baseline) and 27% (AC). PPO is the only agent whose AC performance is meaningfully worse than its baseline: 28% vs. 31%. The annualized Sharpe ratio improves from 0.9 to 1.1 for TD3 when switching from the baseline to AC; other agents remain in the 0.9–1.0 range. TD3 total trading cost decreases by $\sim$15% under AC relative to the baseline, while the remaining agents see $\sim$30% lower costs (Fig. 6).
@@
Figure 6: OOS annualized return—POE, all five agents under 10 bps baseline vs. AC impact.
@@ VII Conclusion
We presented a suite of three Gymnasium-compatible trading environments with realistic, nonlinear market impact models and evaluated five DRL algorithms (A2C, PPO, DDPG, SAC, TD3) across stock trading, margin trading, and portfolio optimization formulations. Our results demonstrate that: (1) the choice of cost model materially affects both absolute performance and the relative ranking of algorithms across all three environments—in stock trading, optimized PPO achieves 20% OOS return under the baseline but only 15% under AC, while TD3 improves from 15% to 18%; in POE, TD3 under AC achieves 32% (best overall) while the same algorithm under baseline yields only 26%; (2) the AC model produces dramatically different trading behavior—non-optimized TD3 in stock trading sees daily costs drop from \$200k to \$8k (96% reduction) when switching from baseline to AC, with turnover from 19% to 1%, while in POE the nonlinear cost model enables genuine training convergence that does not occur under flat fees; (3) hyperparameter optimization is essential not only for OOS performance but for constraining pathological execution—PPO and SAC exhibit rising POV and turnover across epochs without HPO (SAC stock trading costs drop 82% with HPO; PPO POE IS POV reaches 1.39 without HPO vs. 0.36 with); and (4) algorithm choice interacts strongly with both the environment and cost model—DDPG’s margin Sharpe jumps from -2.1 to 0.3 under AC while SAC’s drops from -0.5 to -1.2, PPO is the best stock trading agent under baseline but is hurt by AC, and TD3 benefits most from AC across both stock and POE environments.

### 04 Apr 2026 v2

@@ Abstract
Reinforcement learning (RL) has shown promise for trading, yet most open-source backtesting environments assume negligible or fixed transaction costs, which can cause agents to learn undesirable trading behaviors that fail on real-world execution. We introduce a suite of three Gymnasium-compatible trading environments—MACE (Market-Adjusted Cost Execution) stock trading, margin trading, and portfolio optimization—that integrate nonlinear market impact models grounded in the Almgren–Chriss (AC) framework and the empirically validated square-root impact law. Each environment shares a unified infrastructure: pluggable cost models, permanent impact tracking with exponential decay, and comprehensive trade-level logging. We evaluate five DRL algorithms (A2C, PPO, DDPG, SAC, TD3) on the NASDAQ 100 universe, comparing a fixed 10 bps baseline cost model against the AC model with Optuna-optimized hyperparameters. Our results show that (i) the cost model materially affects both absolute performance and the relative ranking of algorithms across all three environments—in stock trading, optimized PPO achieves the best OOS result (34% total return, Sharpe 1.06) under the baseline but drops to 25% under AC, while TD3 improves from 24% to 30%; in portfolio optimization, TD3 under AC achieves 32% total return (best overall) while the same algorithm under baseline yields only 26% (worst among agents); (ii) the AC model produces dramatically different trading behavior—non-optimized TD3 in stock trading sees daily costs drop from \$200k to \$8k when switching to AC, with turnover falling from 19% to 1%; (iii) hyperparameter optimization (HPO) is essential for constraining pathological trading—without HPO, agents exhibit monotonically increasing participation rates across epochs (e.g., SAC stock trading costs drop 82% with HPO); and (iv) algorithm choice interacts strongly with the cost model in environment-specific ways—in the margin environment, DDPG’s OOS Sharpe jumps from -2.1 to 0.3 under AC while SAC’s drops from -0.5 to -1.2. We release the full environment suite as an open-source extension to FinRL-Meta.
@@ III-C Reward Function
Figure 1: OOS total return—MACE stock trading, all five agents under baseline vs. AC impact, optimized params. Black line: QQEW benchmark (19%).
@@ V-A MACE Stock Trading: Five-Agent Comparison
All configurations outperform the QQEW benchmark (19% OOS total return) except non-optimized TD3 under AC. The best result is optimized PPO under the baseline cost model: 34% total return and 1.06 Sharpe, switching to the AC cost model reduces total return to 25% (Sharpe 1.03), but trading costs drop 55%, average order POV from 0.18 to 0.14, max drawdown from -20% to -16%, turnover from 1.3% to 1.1%, and volatility from 19% to 15%—a substantially less risky portfolio. TD3 shows the opposite pattern: AC improves total returns from 24% to 30% and Sharpe from 0.9 to 1.1, while turnover drops from 4% to 3%, costs fall 13%, and POV decreases from 0.36 to 0.20. DDPG also improves modestly under AC; A2C and SAC are slightly worse.
HPO dramatically reduces trading activity. Optimized SAC cuts turnover from 5% to 2% on AC and trading costs by 82% (POV from 0.46 to 0.18); DDPG shows a similar pattern. All optimized agents place heavy emphasis on cost reduction, sometimes at the expense of raw returns. The most extreme case is non-optimized TD3 under the baseline: average order POV of 1%, daily turnover of 19%, and average daily trading cost of \$200k. The same agent under AC achieves POV of 0.2%, turnover of 1%, and daily cost of \$8k—a 96% reduction that only moderately affects total returns (22% vs. 18%) because lower volatility (17% to 12%) preserves Sharpe (0.8 to 0.9) (Fig. 2).
@@ V-B Margin Trading: Four-Agent Comparison
All agents underperform the QQEW benchmark OOS, although A2C and PPO outperform IS, suggesting overfitting that HPO could mitigate. The AC model’s effect varies significantly by agent. A2C total returns are 5.5% (baseline) and 6.4% (AC) OOS, with IS annualized returns of 13% and 17.5% respectively—the IS–OOS gap is large but the AC model produces slightly better OOS performance. PPO achieves 15% OOS total return under the baseline but only 9% under AC, despite similar turnover ($\sim$3%) and 40% lower trading costs on AC. DDPG and SAC are the weakest performers, and the cost model makes a large difference in opposite directions: DDPG improves dramatically under AC (OOS Sharpe from -2.1 to 0.3, annualized return from -11% to 2%, max drawdown from -23% to -6%) with similar trading costs and POV, while SAC degrades under AC (OOS Sharpe from -0.5 to -1.2, annualized return from -3% to -7%), with lower POV (0.26 to 0.15) but $\sim$10% higher trading costs (Fig. 4).
@@ V-C Portfolio Optimization: Five-Agent Comparison
All agents outperform the benchmark. TD3 exhibits the widest spread across cost models: with the AC model it achieves 32% total return—the highest OOS result of any configuration—while with the 10 bps baseline it returns only 26%, the lowest among all agents. A2C obtains $\sim$30% total return under both models (slightly higher with AC); SAC reaches 30% (baseline) and 29% (AC); DDPG is comparable at 28% (baseline) and 27% (AC). PPO is the only agent whose AC performance is meaningfully worse than its baseline: 28% vs. 31%. The annualized Sharpe ratio improves from 0.9 to 1.1 for TD3 when switching from the baseline to AC; other agents remain in the 0.9–1.0 range. TD3 total trading cost decreases by $\sim$15% under AC relative to the baseline, while the remaining agents see $\sim$30% lower costs (Fig. 6).
@@
Figure 6: OOS total return—POE, all five agents under 10 bps baseline vs. AC impact.
@@ VII Conclusion
We presented a suite of three Gymnasium-compatible trading environments with realistic, nonlinear market impact models and evaluated five DRL algorithms (A2C, PPO, DDPG, SAC, TD3) across stock trading, margin trading, and portfolio optimization formulations. Our results demonstrate that: (1) the choice of cost model materially affects both absolute performance and the relative ranking of algorithms across all three environments—in stock trading, optimized PPO achieves 34% OOS total return under the baseline but only 25% under AC, while TD3 improves from 24% to 30%; in POE, TD3 under AC achieves 32% total return (best overall) while the same algorithm under baseline yields only 26%; (2) the AC model produces dramatically different trading behavior—non-optimized TD3 in stock trading sees daily costs drop from \$200k to \$8k (96% reduction) when switching from baseline to AC, with turnover falling from 19% to 1%, while in POE the nonlinear cost model enables genuine training convergence that does not occur under flat fees; (3) hyperparameter optimization is essential not only for OOS performance but for constraining pathological execution—PPO and SAC exhibit rising POV and turnover across epochs without HPO (SAC stock trading costs drop 82% with HPO; PPO POE IS POV reaches 1.39 without HPO vs. 0.36 with); and (4) algorithm choice interacts strongly with both the environment and cost model—DDPG’s margin Sharpe jumps from -2.1 to 0.3 under AC while SAC’s drops from -0.5 to -1.2, PPO is the best stock trading agent under baseline but is hurt by AC, and TD3 benefits most from AC across both stock and POE environments.